OpenStudy (anonymous):

the product of two consecutive positive even number is 1520. what are the numbers? (enter solutions from smallest to largest)

OpenStudy (anonymous):

38 and 40

OpenStudy (anonymous):

\[n\ (n+2) = 1520\] \[n^2 + 2n -1520 = 0\] Now you could try to factor or use the quadratic equation :) but for this it's much better to guess. Note that: \[\sqrt{1520} = 38.98717737923585\] So 38 and 40 seem really good to try, and easily verified.